# Stabilizing Policy Gradient Methods via Reward Profiling

### v1

@@ 1 Introduction
Reinforcement learning (RL) optimizes an agent’s performance in a stochastic, sequential decision-making task. Policy‐gradient (PG) methods, which directly optimize parameterized policies from sampled trajectories rather than relying on value‐function bootstrapping, form one of the core paradigms in RL (sutton1998reinforcement; schulman2015trust). Its direct formulation hence enables PG’s effective implementations in high‐dimensional and continuous‐control settings, including robotic manipulation (peters2006policy), locomotion (todorov2012mujoco), and autonomous driving (schulman2017proximal), where value‐based approaches are inefficient.
However, PG methods generally suffer from high variance in gradient estimations. Small stochastic fluctuations in early‐episode rewards of the trajectories can propagate through the Monte Carlo estimator (e.g. REINFORCE (williams1992simple)), leading to erratic updates, slow convergence, and occasional collapse of performance (ilyas2018closer; lehmann2024definitiveguidepolicygradients). Classic variance‐reduction techniques, including baseline/advantage normalization and trust‐region constraints, can improve the stability but also bring drawbacks: problem‐specific tuning (chung2021beyond), second‐order solvers (schulman2015trust), or extra computational overhead (wu2017scalable). A unified framework combining sample efficiency, stability, and simplicity that does not rely on problem-specific baselines can be useful. Therefore, a natural question arises:
@@
We perform a simple experiment to entertain this idea. We apply our reward profiling to a simple REINFORCE algorithm, and implement two algorithms under the CartPole environment (brockman2016openai). Figure 1 shows vanilla REINFORCE can suffer from unstable learning and severe performance fluctuations, whereas the reward profiling provides a much more stable, nearly monotonic improvement. This observation shows that the profiling framework can address the fundamental challenge in PG without acquiring any complicated techniques.
@@
Extensive empirical evaluation (§6). We adapt our framework with three representative algorithms: DDPG (lillicrap2016continuous), TRPO (schulman2015trust), and PPO (schulman2017proximal), and design corresponding profiling algorithms. We then evaluate their performance on continuous-control benchmarks. Results show consistent gains in final performance, sample efficiency, and training stability, providing validation for the algorithmic framework. We also port the algorithm to a Unity-ML ”Reacher” DDPG agent with a separate simulation backend to verify that our framework is broadly applicable while yielding faster convergence and lower variance.
@@ 2 Related Work
Algorithm‐centric approaches. Early variance reduction in REINFORCE introduced baselines, which are often a learned value function, to center Monte-Carlo returns, yielding unbiased and lower-variance updates (sutton1999policy). Actor–critic architectures extend this idea by bootstrapping via temporal-difference learning, trading bias for further variance reduction (konda1999actor; sutton1984temporal); however, a misspecified critic step can introduce harmful bias (chen2021closing; olshevsky2023smallgainanalysissingle). As an alternative approach, trust-region methods such as TRPO enforce a KL-constraint to guarantee monotonic policy improvement under certain regularity conditions (schulman2015trust), but rely on second-order solvers, which are expensive. PPO replaces TRPO’s conjugate-gradient and Hessian computations with a clipped surrogate objective, retaining many of the stability benefits while using only first-order updates (schulman2017proximal); nevertheless, it can still exhibit overly conservative steps or exploration failures in complex, contact-rich tasks (wang2019trust). Beyond on-policy methods, off-policy actor-critic algorithms like DDPG (lillicrap2016continuous) and TD3 (fujimoto18addressing) aim for high sample efficiency via replay buffers, but suffer from spiky Q-value estimates and divergent updates without careful regularization (haarnoja2018soft; islam2017reproducibility). Variants such as SAC inject entropy bonuses for exploration and stability (haarnoja2018soft), yet their gains come at the cost of extra hyperparameters and temperature tuning.
@@
Framework-based approaches. An orthogonal strand of work wraps a generic PG optimizer with lightweight checks or corrections. SVRG-style control variates (e.g., SVRPG) freeze a reference policy to reduce gradient variance, at the expense of extra memory and on-policy rollouts (xu2020improved). Momentum injections, such as Nesterov and heavy-ball variants, have been proposed to accelerate PG under smoothness assumptions (xiao2022convergence; chen2024accelerated); although they sometimes amplify early spikes. In acceptance-based schemes, updates are only applied if a held-out performance estimate improves—trace back to optimistic or hysteretic updates in multi-agent settings (omidshafiei2017deep; palma2018multiagent); however, repeatedly validating a ”true” return can be prohibitively costly.
@@
Our reward-profiling wrapper unifies and extends these ideas with three simple, hyperparameter-minimal schemes:Lookback acts like a backtracking line search in policy space, rejecting any update whose empirical returns are lower than the previous policy. Mixup is a convex combination of the old and new parameters to smooth the transition and escape rejection deadlocks. Three-Points evaluates an additional ”midpoint” to choose the best of the old, new, and mixed policies. Requiring only $\mathcal{O}(\epsilon^{-2}\ln(T/\delta))$ extra rollouts, it can be plugged into any PG method (DDPG, TRPO, PPO, etc.) without per-environment tuning. This design delivers theoretical improvement guarantees and consistent empirical gains in final performance, sample efficiency, and training stability. The framework can be related to line‐search and momentum (more1994line; sutskever2013importance; muehlebach2021optimization), but is designed as a virtually zero-hyperparameter, plug‐in wrapper.
@@ 3 Preliminaries
Policy gradient algorithms optimize $J(\theta)$ through gradient ascent, based on the Policy Gradient Theorem (sutton1999policy): $\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\Bigl[\sum_{t=0}^{\infty}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\,Q^{\pi_{\theta}}(s_{t},a_{t})\Bigr]$, where the action‐value function is defined as $Q^{\pi}(s,a)\triangleq\mathbb{E}[G(\tau)|s_{0}=s,a_{0}=a]$.
In practice, as the value functions are unknown, one needs to replace them with the (Monte‐Carlo) estimation, resulting in the straightforward REINFORCE algorithm (sutton1998reinforcement). However, REINFORCE suffers from variances in gradient estimation and unstable performance; hence, modern approaches employ distinct stabilization mechanisms through constrained optimization and off-policy learning. Among on-policy methods, TRPO maximizes a surrogate‐advantage objective subject to a hard KL‐constraint (schulman2015trust):
@@
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\rm old}}(a_{t}\mid s_{t})}$, and $\hat{A}$ being an advantage estimator. TRPO enjoys a theoretical guarantee of (approximate) monotonic policy improvement with the KL constraint. PPO replaces TRPO’s hard constraint with a clipped surrogate that is cheaper to compute, still discourages large updates (schulman2017proximal). Unlike TRPO’s KL‐based trust region, PPO does not admit the same theoretical assurance of monotonic improvement, yet a practical heuristic that works well in large‐scale implementations. Meanwhile, DDPG has remained a standard off-policy PG algorithm, training a deterministic actor $\mu_{\theta}:\mathcal{S}\to\mathcal{A}$ and a critic $Q_{\phi}$ off-policy using a replay buffer $\mathcal{D}$ and slowly-updated target networks (lillicrap2016continuous) $(\mu_{\theta^{\prime}},Q_{\phi^{\prime}})$:
@@
Twin Delayed DDPG (TD3) (fujimoto18addressing) further stabilizes DDPG by (i) using two critics and taking the minimum to reduce overestimation bias, (ii) delaying policy and target network updates, and (iii) adding clipped noise to target actions for policy smoothing.
@@ 4 Reward Profiling Framework
If the noise masks genuine improvements in the noisy estimates, the agent might get stuck. To address this, we draw inspiration from the method mixup, a classic data augmentation technique used in supervised learning (zhang2017mixup), where inputs and labels are convexly combined to reduce inductive bias, widely used in computer vision applications (dutta2024multiview). Specifically, we consider an intermediate policy defined by
$\theta_{\rm mix}=\lambda\,\theta_{\rm new}+(1-\lambda)\,\theta_{\rm old},$
@@
$\theta_{t+1}=\arg\max_{\theta\in\{\theta_{t},\theta_{\rm mix}\}}\hat{J}(\pi_{\theta}).$
We also highlight that, as $\theta_{\rm mix}$ lies closer to $\theta_{t}$ in parameter space (when $\lambda$ is small), this step functions as a cheap trust region, yet without any Hessian computation or KL constraint as in TRPO (schulman2015trust; schulman2017proximal). This step is expected to result in a better performance. Finally, the last modification unifies the best of the first two cases. The Three‐Points variant considers all candidates $\{\theta_{t},\theta^{\prime},\theta_{\rm mix}\}$ and selects the best:
$\theta_{t+1}=\arg\max_{\theta\in\{\theta_{\rm old},\theta_{\rm new},\theta_{\rm mix}\}}\hat{J}(\pi_{\theta}).$
@@ 5 Convergence Analysis
Assumption 1 is standard in policy gradient studies, e.g., (xu2020improved; wang2024non; ganesh2024accelerated). There is a rich family of policies that satisfy the conditions, including the Softmax policy and policies defined through neural networks with smooth and Lipschitz activation functions (tian2023convergence). Finally, we are all set to quote the convergence of our algorithm.
@@
The deterministic PG method (agarwal2021theory; xiao2022convergence) achieves a faster convergence rate than ours because it has access to the true policy gradient without error. In contrast, our algorithm achieves a similar convergence rate when the algorithm is stochastic and the update is based on policy gradient estimation, as in actor-critic algorithms (xu2020improved; suttle2023beyond; olshevsky2023smallgainanalysissingle; chen2021closing; wang2024non). This demonstrates that the convergence rate of our approach is not slowed down much by our profiling framework. More importantly, unlike previous works, Theorem 1 provides a high‐probability guarantee on the last iteration, whereas prior results typically only guarantee the expectation or the best policy found during the learning process. This framework requires additional trajectories of order $\mathcal{O}(\epsilon^{-2}\ln(T/\delta))$, which is still acceptable when the training samples are redundant.
@@
Biased Case. In the above discussions, we consider the case where we compare performances based on Monte-Carlo estimation. In practice, however, it is often the case that PG methods are utilized in actor-critic algorithms, where the value functions are estimated based on the critic. Since the critic part is generally inaccurate, the resulting gradient estimations can have bias. Thus, it is important to extend our framework to the biased case. Given $\theta$, we assume that the critic part provides a biased estimation of the value function $\hat{Q}^{\pi_{\theta}}$, such that $\mathbb{E}[\hat{Q}^{\pi_{\theta}}]={Q}^{\pi_{\theta}}+\epsilon_{\theta},$ where $\epsilon_{\theta}$ is the bias introduced by the errors of the critic part. Such results can be obtained, for instance, when a deep neural network is used to estimate value functions, e.g., (du2019gradient; neyshabur2017implicit; miyato2018spectral), and is widely adapted in actor-critic analysis, e.g., (wang2019neural; zhou2023single; chen2023global; zhang2020provably; qiu2021finite; kumar2023sample; xu2020non; xu2020improving; suttle2023beyond). It can be further shown that when the value function estimation is biased, the resulting gradient estimation has a bias of $CA\epsilon_{\theta}$, as long as $\|\nabla\pi_{\theta}\|\leq C$ (sutton1999policy) and $A=|\mathcal{A}|$ actions. Therefore, without loss of generality, we assume the bias of zero-th and first order gradient of $Q^{\pi_{\theta}}$ are both bounded by some $\epsilon_{\theta}$. We then show the convergence of 1, when the Lookback technique is applied to REINFORCE.
@@
The result shows that our algorithm converges to a stationary point, as long as the value estimation is accurate enough. Moreover, due to the gradient dominance property of the value function (agarwal2021theory), our result also implies the global optimality of the learned policy.
Figure 2: Average Return vs. timesteps (up to 100K) for 3 policy‐gradient methods, PPO (top row), TRPO (middle row), and DDPG (bottom row), on 4 continuous‐control benchmarks: LunarLanderContinuous, Ant, Hopper, Walker2D, and under variants:   Vanilla,   Lookback,   Mixup,   Three‐Points, with $E=5$ for minimal overhead. Extra rollouts tuning gets convergence gains in complex environments.
@@ 6.1 Simulated Gymnasium Environments
We benchmarked our reward-profiling wrapper by performing extensive experiments to stress-test policy learning across the major flavors of policy-gradient methods. We wrapped three canonical algorithms. TRPO, the trusted ”second-order” method with hard KL-constraints, PPO, its ”first-order” surrogate-clipping successor and current on-policy workhorse, and DDPG, an off-policy, replay buffer actor-critic known for sample-efficient continuous control but notoriously unstable, in our profiling layer.
Profiling framework integration. We built a modular profiling wrapper atop Stable-Baselines3 (v1.9) and SB3-Contrib (stable-baselines3), leveraging Gymnasium and PyBullet (benelot2018; towers2024gymnasiumstandardinterfacereinforcement) for simulation. Each variant ran for 100k environment steps across five random seeds. Each profiling round consumes the same 1010k environment steps as the baseline; samples used for improvement checks are reused for training updates. Thus, no additional interactions are collected. All variants share the same per-round evaluation schedule and count. Every candidate set is scored with short episodes ($E=5$), and averages are computed on the identical rollouts for all baselines for fair comparison. The policy network in SB3 is actor-critic, which was used throughout the experiments.
@@
Figure 3: Multi-agent Reacher task in UnityML: multiple robotic arms must coordinate to reach and manipulate randomly placed targets (spheres) on a raised platform, testing control under interaction.
Figure 4: Average return over training in the UnityML Reacher environment, showing profiling-enhanced DDPG not only converges faster but also maintains greater stability compared to its vanilla counterpart.

### v2

@@ 1 Introduction
Reinforcement learning (RL) optimizes an agent’s performance in a stochastic, sequential decision-making task. Policy‐gradient (PG) methods, which directly optimize parameterized policies from sampled trajectories rather than relying on value‐function bootstrapping, form one of the core paradigms in RL (Sutton et al. 1998; Schulman et al. 2015). Its direct formulation hence enables PG’s effective implementations in high‐dimensional and continuous‐control settings, including robotic manipulation (Peters and Schaal 2006), locomotion (Todorov et al. 2012), and autonomous driving (Schulman et al. 2017), where value‐based approaches are inefficient.
However, PG methods generally suffer from high variance in gradient estimations. Small stochastic fluctuations in early‐episode rewards of the trajectories can propagate through the Monte Carlo estimator (e.g. REINFORCE (Williams 1992)), leading to erratic updates, slow convergence, and occasional collapse of performance (Ilyas et al. 2018; Lehmann 2024). Classic variance‐reduction techniques, including baseline/advantage normalization and trust‐region constraints, can improve the stability but also bring drawbacks: problem‐specific tuning (Chung et al. 2021), second‐order solvers (Schulman et al. 2015), or extra computational overhead (Wu et al. 2017). A unified framework combining sample efficiency, stability, and simplicity that does not rely on problem-specific baselines can be useful. Therefore, a natural question arises:
@@
We perform a simple experiment to entertain this idea. We apply our reward profiling to a simple REINFORCE algorithm, and implement two algorithms under the CartPole environment (Brockman et al. 2016). Figure 1 shows vanilla REINFORCE can suffer from unstable learning and severe performance fluctuations, whereas the reward profiling provides a much more stable, nearly monotonic improvement. This observation shows that the profiling framework can address the fundamental challenge in PG without acquiring any complicated techniques.
@@
Extensive empirical evaluation (§6). We adapt our framework with three representative algorithms: DDPG (Lillicrap et al. 2015), TRPO (Schulman et al. 2015), and PPO (Schulman et al. 2017), and design corresponding profiling algorithms. We then evaluate their performance on continuous-control benchmarks. Results show consistent gains in final performance, sample efficiency, and training stability, providing validation for the algorithmic framework. We also port the algorithm to a Unity-ML “Reacher” DDPG agent with a separate simulation backend to verify that our framework is broadly applicable while yielding faster convergence and lower variance.
@@ 2 Related Work
Algorithm‐centric approaches. Early variance reduction in REINFORCE introduced baselines, which are often a learned value function, to center Monte-Carlo returns, yielding unbiased and lower-variance updates (Sutton et al. 1999). Actor–critic architectures extend this idea by bootstrapping via temporal-difference learning, trading bias for further variance reduction (Konda and Tsitsiklis 1999; Sutton 1984); however, a misspecified critic step can introduce harmful bias (Chen et al. 2021; Olshevsky and Gharesifard 2023). As an alternative approach, trust-region methods such as TRPO enforce a KL-constraint to guarantee monotonic policy improvement under certain regularity conditions (Schulman et al. 2015), but rely on second-order solvers, which are expensive. PPO replaces TRPO’s conjugate-gradient and Hessian computations with a clipped surrogate objective, retaining many of the stability benefits while using only first-order updates (Schulman et al. 2017); nevertheless, it can still exhibit overly conservative steps or exploration failures in complex, contact-rich tasks (Wang et al. 2019b). Beyond on-policy methods, off-policy actor-critic algorithms like DDPG (Lillicrap et al. 2015) and TD3 (Fujimoto et al. 2018) aim for high sample efficiency via replay buffers, but suffer from spiky Q-value estimates and divergent updates without careful regularization (Haarnoja et al. 2018; Islam et al. 2017). Variants such as SAC inject entropy bonuses for exploration and stability (Haarnoja et al. 2018), yet their gains come at the cost of extra hyperparameters and temperature tuning.
@@
Framework-based approaches. An orthogonal strand of work wraps a generic PG optimizer with lightweight checks or corrections. SVRG-style control variates (e.g., SVRPG) freeze a reference policy to reduce gradient variance, at the expense of extra memory and on-policy rollouts (Xu et al. 2018). Momentum injections, such as Nesterov and heavy-ball variants, have been proposed to accelerate PG under smoothness assumptions (Xiao 2022; Chen et al. 2024); although they sometimes amplify early spikes. In acceptance-based schemes, updates are only applied if a held-out performance estimate improves—trace back to optimistic or hysteretic updates in multi-agent settings (Omidshafiei et al. 2017; Palma et al. 2018); however, repeatedly validating a “true” return can be prohibitively costly.
@@
Our reward-profiling wrapper unifies and extends these ideas with three simple, hyperparameter-minimal schemes:Lookback acts like a backtracking line search in policy space, rejecting any update whose empirical returns are lower than the previous policy. Mixup is a convex combination of the old and new parameters to smooth the transition and escape rejection deadlocks. Three-Points evaluates an additional “midpoint” to choose the best of the old, new, and mixed policies. Requiring only $\mathcal{O}(\epsilon^{-2}\ln(T/\delta))$ extra rollouts, it can be plugged into any PG method (DDPG, TRPO, PPO, etc.) without per-environment tuning. This design delivers theoretical improvement guarantees and consistent empirical gains in final performance, sample efficiency, and training stability. The framework can be related to line‐search and momentum (Moré et al. 1994; Sutskever et al. 2013; Muehlebach and Jordan 2021), but is designed as a virtually zero-hyperparameter, plug‐in wrapper.
@@ 3 Preliminaries
Policy gradient algorithms optimize $J(\theta)$ through gradient ascent, based on the Policy Gradient Theorem (Sutton et al. 1999): $\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\Bigl[\sum_{t=0}^{\infty}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\,Q^{\pi_{\theta}}(s_{t},a_{t})\Bigr]$, where the action‐value function is defined as $Q^{\pi}(s,a)\triangleq\mathbb{E}[G(\tau)|s_{0}=s,a_{0}=a]$.
In practice, as the value functions are unknown, one needs to replace them with the (Monte‐Carlo) estimation, resulting in the straightforward REINFORCE algorithm (Sutton et al. 1998). However, REINFORCE suffers from variances in gradient estimation and unstable performance; hence, modern approaches employ distinct stabilization mechanisms through constrained optimization and off-policy learning. Among on-policy methods, TRPO maximizes a surrogate‐advantage objective subject to a hard KL‐constraint (Schulman et al. 2015):
@@
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\rm old}}(a_{t}\mid s_{t})}$, and $\hat{A}$ being an advantage estimator. TRPO enjoys a theoretical guarantee of (approximate) monotonic policy improvement with the KL constraint. PPO replaces TRPO’s hard constraint with a clipped surrogate that is cheaper to compute, still discourages large updates (Schulman et al. 2017). Unlike TRPO’s KL‐based trust region, PPO does not admit the same theoretical assurance of monotonic improvement, yet a practical heuristic that works well in large‐scale implementations. Meanwhile, DDPG has remained a standard off-policy PG algorithm, training a deterministic actor $\mu_{\theta}:\mathcal{S}\to\mathcal{A}$ and a critic $Q_{\phi}$ off-policy using a replay buffer $\mathcal{D}$ and slowly-updated target networks (Lillicrap et al. 2015) $(\mu_{\theta^{\prime}},Q_{\phi^{\prime}})$:
@@
Twin Delayed DDPG (TD3) (Fujimoto et al. 2018) further stabilizes DDPG by (i) using two critics and taking the minimum to reduce overestimation bias, (ii) delaying policy and target network updates, and (iii) adding clipped noise to target actions for policy smoothing.
@@ 4 Reward Profiling Framework
If the noise masks genuine improvements in the noisy estimates, the agent might get stuck. To address this, we draw inspiration from the method mixup, a classic data augmentation technique used in supervised learning (Zhang et al. 2017), where inputs and labels are convexly combined to reduce inductive bias, widely used in computer vision applications (Dutta et al. 2024). Specifically, we consider an intermediate policy defined by
$\theta_{\rm mix}=\lambda\,\theta_{\rm new}+(1-\lambda)\,\theta_{\rm old},$ (4)
@@
$\theta_{t+1}=\arg\max_{\theta\in\{\theta_{t},\theta_{\rm mix}\}}\hat{J}(\pi_{\theta}).$ (5)
We also highlight that, as $\theta_{\rm mix}$ lies closer to $\theta_{t}$ in parameter space (when $\lambda$ is small), this step functions as a cheap trust region, yet without any Hessian computation or KL constraint as in TRPO (Schulman et al. 2015, 2017). This step is expected to result in a better performance. Finally, the last modification unifies the best of the first two cases. The Three‐Points variant considers all candidates $\{\theta_{t},\theta^{\prime},\theta_{\rm mix}\}$ and selects the best:
$\theta_{t+1}=\arg\max_{\theta\in\{\theta_{\rm old},\theta_{\rm new},\theta_{\rm mix}\}}\hat{J}(\pi_{\theta}).$ (6)
@@ 5 Convergence Analysis
Assumption 1 is standard in policy gradient studies, e.g., (Xu et al. 2018; Wang et al. 2024; Ganesh and Aggarwal 2024). There is a rich family of policies that satisfy the conditions, including the Softmax policy and policies defined through neural networks with smooth and Lipschitz activation functions (Tian et al. 2023). Finally, we are all set to quote the convergence of our algorithm.
@@
The deterministic PG method (Agarwal et al. 2021; Xiao 2022) achieves a faster convergence rate than ours because it has access to the true policy gradient without error. In contrast, our algorithm achieves a similar convergence rate when the algorithm is stochastic and the update is based on policy gradient estimation, as in actor-critic algorithms (Xu et al. 2018; Suttle et al. 2023; Olshevsky and Gharesifard 2023; Chen et al. 2021; Wang et al. 2024). This demonstrates that the convergence rate of our approach is not slowed down much by our profiling framework. More importantly, unlike previous works, Theorem 1 provides a high‐probability guarantee on the last iteration, whereas prior results typically only guarantee the expectation or the best policy found during the learning process. This framework requires additional trajectories of order $\mathcal{O}(\epsilon^{-2}\ln(T/\delta))$, which is still acceptable when the training samples are redundant.
@@
Biased Case. In the above discussions, we consider the case where we compare performances based on Monte-Carlo estimation. In practice, however, it is often the case that PG methods are utilized in actor-critic algorithms, where the value functions are estimated based on the critic. Since the critic part is generally inaccurate, the resulting gradient estimations can have bias. Thus, it is important to extend our framework to the biased case. Given $\theta$, we assume that the critic part provides a biased estimation of the value function $\hat{Q}^{\pi_{\theta}}$, such that $\mathbb{E}[\hat{Q}^{\pi_{\theta}}]={Q}^{\pi_{\theta}}+\epsilon_{\theta},$ where $\epsilon_{\theta}$ is the bias introduced by the errors of the critic part. Such results can be obtained, for instance, when a deep neural network is used to estimate value functions, e.g., (Du et al. 2019; Neyshabur 2017; Miyato et al. 2018), and is widely adapted in actor-critic analysis, e.g., (Wang et al. 2019a; Zhou and Lu 2023; Chen et al. 2023; Zhang et al. 2020; Qiu et al. 2021; Kumar et al. 2023; Xu et al. 2020b, a; Suttle et al. 2023). It can be further shown that when the value function estimation is biased, the resulting gradient estimation has a bias of $CA\epsilon_{\theta}$, as long as $\|\nabla\pi_{\theta}\|\leq C$ (Sutton et al. 1999) and $A=|\mathcal{A}|$ actions. Therefore, without loss of generality, we assume the bias of zero-th and first order gradient of $Q^{\pi_{\theta}}$ are both bounded by some $\epsilon_{\theta}$. We then show the convergence of 1, when the Lookback technique is applied to REINFORCE.
@@
The result shows that our algorithm converges to a stationary point, as long as the value estimation is accurate enough. Moreover, due to the gradient dominance property of the value function (Agarwal et al. 2021), our result also implies the global optimality of the learned policy.
Figure 2: Average Return vs. timesteps ( $\times 10^{4}$) for 3 policy-gradient methods, PPO (top row), TRPO (middle row), and DDPG (bottom row), across 4 continuous-control benchmarks: LunarLanderCont, Hopper, Walker2D, Ant under variants: —— Vanilla, – – – Lookback, –·–·– Mixup, ····· Three-Points, with $E=5$ for minimal overhead. Shaded regions indicate $\pm 1$ standard deviation. Extra rollouts tuning yields convergence gains in complex environments.
@@ 6.1 Simulated Gymnasium Environments
We benchmarked our reward-profiling wrapper by performing extensive experiments to stress-test policy learning across the major flavors of policy-gradient methods. We wrapped three canonical algorithms. TRPO, the trusted “second-order” method with hard KL-constraints, PPO, its “first-order” surrogate-clipping successor and current on-policy workhorse, and DDPG, an off-policy, replay buffer actor-critic known for sample-efficient continuous control but notoriously unstable, in our profiling layer.
Profiling framework integration. We built a modular profiling wrapper atop Stable-Baselines3 (v1.9) and SB3-Contrib (Raffin et al. 2021), leveraging Gymnasium and PyBullet (Ellenberger 2018; Towers et al. 2024) for simulation. Each variant ran for 100k environment steps across five random seeds. Each profiling round consumes the same 1010k environment steps as the baseline; samples used for improvement checks are reused for training updates. Thus, no additional interactions are collected. All variants share the same per-round evaluation schedule and count. Every candidate set is scored with short episodes ($E=5$), and averages are computed on the identical rollouts for all baselines for fair comparison. The policy network in SB3 is actor-critic, which was used throughout the experiments.
@@
Figure 3: Multi-agent Reacher task in UnityML: multiple arms coordinate to reach and manipulate randomly placed targets on a raised platform, testing control under interaction.
Figure 4: Perormance in the UnityML Reacher environment. Profiling-enhanced DDPG converges faster and maintains greater stability than the vanilla counterpart.